# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 1.2 $e$-Newton number and Algebraic degrees
(see Corollary 5.49) The polar degree of a homogeneous Newton non-degenerate polynomial is bounded by the local $h^{*}$-polynomial of the Cayley sum of its Newton polytope with the unit simplex.
We also discuss the important cases of polar degree equal to zero and one (corresponding to vanishing hessian and homaloidal hypersurfaces), see Remark 5.45, Question 5.46 and Conjecture 5.47. Similar questions are interesting for the maximum likelihood degree, see Question 5.40.
@@ 2 Preliminaries: Katz-Stapledon decomposition formulas
In this section we recall (lower) Eulerian posets that abstract face posets of polytopes and their $h$- and $g$-polynomials. We recall $h^{*}$- and local $h^{*}$-polynomials of polytopes coming from the Ehrhart theory. We also recall the local $h$-polynomials of strong formal subdivisions and Katz-Stapledon decomposition formulas from [KS16].
For an introduction to posets (partially ordered sets) see [Stan]. For surveys on the local $h$-polynomial of triangulations of simplices see [Cha] and [Ath]. For a survey on the local $h$-polynomials of polyhedral subdivisions and Katz-Stapledon decomposition formulas see [BKN24, §2]. For an introduction to the Ehrhart theory including $h^{*}$-polynomial see [BR07]. For a survey on local $h^{*}$-polynomials see [BKN24, §2].
2.1 Eulerian posets, $h$- and $g$-polynomial
@@
The dual of a finite poset $\mathcal{P}$ is denoted $\mathcal{P}^{\ast}$. A finite poset $\mathcal{P}$ is locally graded if every inclusion-maximal chain in every interval $[x,y]$ has the same length $r(x,y)$. The rank $\operatorname{rk}(\mathcal{P})$ is the length of the longest chain in $\mathcal{P}$. If in addition there exists a rank function $\rho:\mathcal{P}\rightarrow\mathbb{Z}$, i.e., $r(x,y)=\rho(y)-\rho(x)$ for every interval $[x,y]$, then $\mathcal{P}$ is called ranked. If $\mathcal{P}$ is ranked and every interval $[x,y]$ with $x\neq y$ has the same number of even rank and odd rank elements, then $\mathcal{P}$ is locally Eulerian. If $\mathcal{P}$ is locally Eulerian and contains a minimal element $\hat{0}$, then it is called lower Eulerian. If it also contains a maximum $\hat{1}$, then $\mathcal{P}$ is called Eulerian. In presence of a minimum $\hat{0}$ in a ranked poset $\mathcal{P}$, we will always assume that the rank function satisfies $\rho(\hat{0})=0$.
@@
We use the notations of $g$- and $h$-polynomials from [KS16]. Unfortunately, these notations are a bit different in different literature. For more details, see [BKN24, Remark 2.6.].
@@
Let $B$ be an Eulerian poset of rank $n$. If $n=0$, then $g(B;t)=1$. If $n>0$, then $g(B;t)\in\mathbb{Z}[t]$ is the unique polynomial in $t$ of degree strictly less the $n/2$ satisfying
@@
(see e.g. [KS16, Example 3.8.]). The constant term of $g(B;t)$ is $g(B;0)=1$. The linear coefficient of $g(B;t)$ is #$\#\{x\in B|\rho(\hat{0}_{B},x)=1\}-n$.
@@
(see e.g. [KS16, Example 3.9.]) One can verify that $g(B;t)=1$ if and only if $B$ is a Boolean algebra.
@@
(see e.g. [KS16, Example 3.13]) Let $B$ be a lower Eulerian poset of rank $n$. Then the constant term of $h(B;t)$ is $h(B;0)=1$, and the linear coefficient of $h(B;t)$ is #$\#\{x\in B|\rho(\hat{0}_{B},x)=1\}-n$.
Claim 2.8.
If $B$ is an Eulerian poset of rank $n$, then comparison of the recursive formulas in Definition 2.3 and Definition 2.6 implies that $h(B;t)=g(B;t)$. If, furthermore, $n>0$, then $B\setminus\{\hat{1}\}$ is a lower Eulerian poset of rank $n-1$, and
@@
In particular, $h(B\setminus\{\hat{1}\};t)$ is a polynomial of degree $n-1$ with symmetric coefficients.
@@
The poset of faces of a polytope $P$ (including the empty face) is an Eulerian poset under inclusion, called the face poset $\mathcal{P}$ of $P$, with $\rho(Q)=\dim Q+1$ for every face $Q$ of $P$. The poset of faces of a cone $C$ (excluding the empty face) is an Eulerian poset under inclusion, called the face poset $\mathcal{C}$ of $C$, with $\rho(F)=\dim F$ for every face $F$ of $C$. If the cone $C$ is sharp then its face poset coincides with the face poset of the polytope in its hyperplane section.
@@
2.2 Local $h$-polynomials of strong formal subdivisions
Strong formal subdivisions abstract polyhedral subdivisions. In this subsection we recall the Katz-Stapledon decomposition formulas for the local $h$-polynomial of strong formal subdivisions from [KS16].
A function $f:\Gamma\rightarrow B$ between finite posets is order-preserving if $y\leq y^{\prime}\in\Gamma$ implies $f(y)\leq f(y^{\prime})\in B$. If $(\Gamma,r_{\Gamma})$ and $(B,r_{B})$ are ranked posets, then a function $f:\Gamma\rightarrow B$ is rank-increasing if $r_{\Gamma}(y)\leq r_{B}(f(y))$ for all $y\in\Gamma$. Throughout, we let $\sigma:\Gamma\rightarrow B$ be an order-preserving, rank-increasing function between locally Eulerian posets with rank functions $\rho_{\Gamma}$ and $\rho_{B}$ respectively.
@@ 2.2 Local $h$-polynomials of strong formal subdivisions
The following definition abstracts polyhedral subdivisions. The property of subdivisions that is abstracted is that the subdivision of the relative interior of a polytope has Euler characteristic equal to $1$.
@@
If $\Gamma$ and $B$ are lower Eulerian, then it follows that $\sigma(\hat{0}_{\Gamma})=\hat{0}_{B}$, and the rank $\operatorname{rk}(\sigma)\in\mathbb{Z}_{\geq 0}$ of $\sigma$ is defined to be $\operatorname{rk}(\sigma):=\rho_{B}(\hat{0}_{B})-\rho_{\Gamma}(\hat{0}_{\Gamma})$. Alternatively, it follows from the fact that $\sigma$ is rank-increasing and strongly surjective that $\operatorname{rk}(\sigma)=\operatorname{rk}(\Gamma)-\operatorname{rk}(B)$.
@@
[KS16, Definition 4.1] Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ and an Eulerian poset $B$. Then the local $\mathbf{h}$-polynomial $l_{B}(\Gamma;t)\in\mathbb{Z}[t]$ is defined by
$l_{B}(\Gamma;t)=\sum_{x\in B}h(\Gamma_{x};t)(-1)^{\rho_{B}(x,\hat{1}_{B})}g([x,\hat{1}_{B}]^{*};t).$
We also introduce the following useful notation. Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between locally Eulerian posets with rank functions $\rho_{\Gamma}$ and $\rho_{B}$ respectively, and fix $y\in\Gamma$ and $x\in B$. If $\sigma(y)\leq x$, then we may the restricted strong formal subdivision $\sigma:(\Gamma_{\geq y})_{x}\rightarrow[\sigma(y),x]$. In this case we set
$l_{B}(\Gamma,x,y;t):=l_{[\sigma(y),x]}((\Gamma_{\geq y})_{x};t)=\sum_{\sigma(y)\leq x^{\prime}\leq x}h((\Gamma_{\geq y})_{x^{\prime}};t)(-1)^{\rho_{B}(x^{\prime},x)}g([x^{\prime},x]^{*};t).$
We set $l_{B}(\Gamma,x,y;t)=0$ if $\sigma(y)\nleq x$.
@@
Corollary 2.16.
[KS16, Corollary 4.5.] Let $\sigma:\Gamma\to B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ of rank $r$ and an Eulerian poset $B$. Then the local $h$-polynomial $\ell_{B}(\Gamma;t)$ has symmetric coefficients centered at $r/2$.
@@
Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ and an Eulerian poset $B$. We recover the $h$-polynomial of $\Gamma$ via:
@@
Corollary 2.18.
(Katz-Stapledon decomposition of local $h$-polynomial)[KS16, Corollary 4.7.] Let $\tau:\Omega\rightarrow\Gamma$ and let $\sigma:\Gamma\rightarrow B$ be strong formal subdivisions lower Eulerian posets $\Omega,\Gamma$ and Eulerian poset $B$. Then for $z\in\Omega$, $x\in\Gamma$,
@@
The Boolean algebra on $r$ elements consists of all subsets of a set of cardinality $r$ and forms an Eulerian poset under inclusion of rank $r$. It is the face poset of an $(r-1)$-dimensional simplex. A poset $B$ with minimal element $\hat{0}$ is simplicial if for every $x$ in $B$, the interval $[\hat{0},x]$ is a Boolean algebra.
@@
The local $h$-polynomial of strong formal subdivision is not necessarily non-negative (see [KS16, Example 5.6.] and Corollary 3.12).
In case of polyhedral subdivision $\mathcal{S}$ of a polytope $P$ we denote $\ell_{P}(\mathcal{S};t)$ assuming the strong formal subdivision between the corresponding face posets. Recall that $\sigma(F),F\in\mathcal{S}$ is defined as the minimal face of $P$ containing $F$. For any face $Q$ of $P$ containing $\sigma(F)$, we may consider the corresponding local $h$-polynomial:
$l_{P}(\mathcal{S},Q,F;t):=l_{[\emptyset,P]}(\mathcal{S},Q,F;t)=l_{[\sigma(F),Q]}(\operatorname{St}_{\mathcal{S}|_{Q}}(F);t).$
@@
[KS16, Theorem 6.1.] Let $\mathcal{S}$ be a polyhedral subdivision of a polytope P. Then for every cell $F$ of $\mathcal{S}$ and face $Q$ of $P$ containing $F$, the local $h$-polynomial $\ell_{P}(\mathcal{S},Q,F;t)$ has non-negative symmetric coefficients. Moreover, if $\mathcal{S}$ is a regular subdivision, then the coefficients of $\ell_{P}(\mathcal{S},Q,F;t)$ are unimodal.
@@
In [KS16, Theorem 6.1.] the condition ‘‘symmetric’’ is written only for regular subdivisions. But due to [KS16, Corollary 4.5] the local $h$-polynomial has symmetric coefficients for arbitrary strong formal subdivision, therefore we include the condition ‘‘symmetric’’ for arbitrary polyhedral subdivision. We also removed the word ‘‘rational’’ because the results from [Kar19] allows this (see e.g. [BKN24, Theorem 2.11]). The latter remark is non-trivial since there are non-rational convex polytopes which are not combinatorially equivalent to any rational polytopes, see e.g. [Zie08].
The local $h$-polynomials are non-negative in the polyhedral case because they appear when applying the Beilinson-Bernstein-Deligne decomposition theorem from [BBD81] to the toric morphisms (see [KS16, Lemma 6.4] and [CMM18]; see also [GKZ94, Notes and References, Part 2, Chapter 11; p. 514]).
2.3 $h^{*}$- and local $h^{*}$-polynomials of lattice polytopes
In this subsection we recall Ehrhart theory, the (local) $h^{*}$-polynomials and Katz-Stapledon decomposition formulas. We recommend [BR07] as an introduction to Ehrhart theory.
Let $P$ be a non-empty lattice polytope in a lattice $\mathbb{Z}^{d}$. Consider the function $f_{P}(m)=\#(mP\cap\mathbb{Z}^{d})$, for $m\in\mathbb{Z}_{>0}$. By Ehrhart’s theorem [BR07, Section 3.3], $f_{P}(m)$ is a polynomial of degree $\dim P$, called the Ehrhart polynomial of $P$, and $f_{P}(0)=1$. It follows that we can write

### v2

@@ 1.2 $\mathbf{e}$-Newton number and Algebraic degrees
(see Corollary 5.56.) The polar degree of a homogeneous Newton non-degenerate polynomial is bounded by the local $h^{*}$-polynomial of the Cayley sum of its Newton polytope with the unit simplex.
We also discuss the important cases of polar degree equal to zero and one (corresponding to vanishing hessian and homaloidal hypersurfaces), see Remark 5.52, Question 5.53 and Conjecture 5.54. Similar questions are interesting for the maximum likelihood degree (see Question 5.46).
@@ 2 Preliminaries: Katz-Stapledon decomposition formulas
In this section we recall (lower) Eulerian posets that abstract face posets of polytopes and their $h$- and $g$-polynomials. We recall $h^{*}$- and local $h^{*}$-polynomials of polytopes defined in [Stan92] coming from Ehrhart theory. We also recall the local $h$-polynomials of strong formal subdivisions and Katz-Stapledon decomposition formulas from [KS16].
For an introduction to posets (partially ordered sets), $h$- and $g$-polynomials, see [Stan]. For surveys on the local $h$-polynomial of triangulations of simplices see [Cha] and [Ath]. For a survey on the local $h$-polynomials of polyhedral subdivisions and Katz-Stapledon decomposition formulas see [BKN24, §2]. For an introduction to the Ehrhart theory including $h^{*}$-polynomial see [BR07]. For a survey on local $h^{*}$-polynomials see [BKN24, §2].
2.1 Eulerian posets, $\mathbf{h}$- and $\mathbf{g}$-polynomial
Eulerian posets were defined by Stanley in [Stan81]. For instance, they abstract face posets of polytopes (see Example 2.9).
@@ 2.1 Eulerian posets, $\mathbf{h}$- and $\mathbf{g}$-polynomial
The dual of a finite poset $\mathcal{P}$ is denoted $\mathcal{P}^{\ast}$. A finite poset $\mathcal{P}$ is locally graded if every inclusion-maximal chain in every interval $[x,y]$ has the same length $r(x,y)$. The rank $\operatorname{rk}(\mathcal{P})$ is the length of the longest chain in $\mathcal{P}$. If in addition there exists a rank function $\rho:\mathcal{P}\rightarrow\mathbb{Z}$, i.e., $r(x,y)=\rho(y)-\rho(x)$ for every interval $[x,y]$, then $\mathcal{P}$ is called ranked. If $\mathcal{P}$ is ranked and every interval $[x,y]$ with $x\neq y$ has the same number of even rank and odd rank elements, then $\mathcal{P}$ is locally Eulerian. If $\mathcal{P}$ is locally Eulerian and contains a minimal element $\hat{0}$, then it is called lower Eulerian. If it also contains a maximum $\hat{1}$, then $\mathcal{P}$ is called Eulerian. In the presence of a minimum $\hat{0}$ in a ranked poset $\mathcal{P}$, we will always assume that the rank function satisfies $\rho(\hat{0})=0$.
@@
We use the notations of $g$- and $h$-polynomials from [KS16]. Unfortunately, these notations are a bit different in different literature. For more details, see [BKN24, Remark 2.6].
@@
Let $B$ be an Eulerian poset of rank $n$. If $n=0$, then $g(B;t)=1$. If $n>0$, then $g(B;t)\in\mathbb{Z}[t]$ is the unique polynomial in $t$ of degree strictly less than $n/2$ satisfying
@@
The constant term of $g(B;t)$ is $g(B;0)=1$. The linear coefficient of $g(B;t)$ is $\{x\in B|\rho(\hat{0}_{B},x)=1\}-n$.
@@
[BN08, Remark 4.2] One can verify that $g(B;t)=1$ if and only if $B$ is a Boolean algebra.
@@
Let $B$ be a lower Eulerian poset of rank $n$. Then the constant term of $h(B;t)$ is $h(B;0)=1$, and the linear coefficient of $h(B;t)$ is $\{x\in B|\rho(\hat{0}_{B},x)=1\}-n$.
Remark 2.8.
(see e.g. [KS16, Example 3.14]) If $B$ is an Eulerian poset of rank $n$, then comparison of the recursive formulas in Definition 2.3 and Definition 2.6 implies that $h(B;t)=g(B;t)$. If, furthermore, $n>0$, then $B\setminus\{\hat{1}\}$ is a lower Eulerian poset of rank n1n-1, and
@@
In particular, $h(B\setminus\{\hat{1}\};t)$ is a polynomial of degree n1n-1 with symmetric coefficients.
@@
The poset of faces of a polytope $P$ (including the empty face) is an Eulerian poset under inclusion, called the face poset $\mathcal{P}$ of $P$, with $\rho(Q)=\dim Q+1$ for every face $Q$ of $P$. The poset of faces of a cone $C$ (excluding the empty face) is an Eulerian poset under inclusion, called the face poset $\mathcal{C}$ of $C$, with $\rho(F)=\dim F$ for every face $F$ of $C$. If the cone $C$ is sharp (i.e., pointed), then its face poset is isomorphic to the face poset of any of its compact cross-sectional polytopes.
@@
2.2 Local $\mathbf{h}$-polynomials of strong formal subdivisions
Strong formal subdivisions abstract polyhedral subdivisions. In this subsection we recall Katz-Stapledon decomposition formulas for the local $h$-polynomial of strong formal subdivisions from [KS16]. These definitions refine Stanley’s definitions of formal subdivision and local $h$-polynomial from [Stan92, Definition 7.4] and [Stan92, Corollary 7.7]; see [KS16, Remark 3.33]. The local $h$-polynomials were defined by Stanley in [Stan92] in order to prove the monotonicity of the $h$-vector under subdivisions of complexes.
A function $f:\Gamma\rightarrow B$ between finite posets is order-preserving if $y\leq y^{\prime}\in\Gamma$ implies $f(y)\leq f(y^{\prime})\in B$. If $(\Gamma,r_{\Gamma})$ and $(B,r_{B})$ are ranked posets, then a function $f:\Gamma\rightarrow B$ is rank-increasing if $r_{\Gamma}(y)\leq r_{B}(f(y))$ for all $y\in\Gamma$. Throughout, we let $\sigma:\Gamma\rightarrow B$ be an order-preserving, rank-increasing function between locally Eulerian posets with rank functions $\rho_{\Gamma}$ and $\rho_{B}$, respectively.
@@ 2.2 Local $\mathbf{h}$-polynomials of strong formal subdivisions
The following definition abstracts polyhedral subdivisions.
@@
If $\Gamma$ and $B$ are lower Eulerian, then it follows that $\sigma(\hat{0}_{\Gamma})=\hat{0}_{B}$, and the rank $\operatorname{rk}(\sigma)\in\mathbb{Z}_{\geq 0}$ of $\sigma$ is defined as $\operatorname{rk}(\sigma):=\rho_{B}(\hat{0}_{B})-\rho_{\Gamma}(\hat{0}_{\Gamma})$. Alternatively, it follows from the fact that $\sigma$ is rank-increasing and strongly surjective that $\operatorname{rk}(\sigma)=\operatorname{rk}(\Gamma)-\operatorname{rk}(B)$.
@@
Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ and an Eulerian poset $B$. Then the local $\mathbf{h}$-polynomial $\ell_{B}(\Gamma;t)\in\mathbb{Z}[t]$ is defined by
$\ell_{B}(\Gamma;t)=\sum_{x\in B}h(\Gamma_{x};t)(-1)^{\rho_{B}(x,\hat{1}_{B})}g([x,\hat{1}_{B}]^{*};t).$
We also introduce the following useful notation. Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between locally Eulerian posets with rank functions $\rho_{\Gamma}$ and $\rho_{B}$ respectively, and fix $y\in\Gamma$ and $x\in B$. If $\sigma(y)\leq x$, then we may consider the restricted strong formal subdivision $\sigma:(\Gamma_{\geq y})_{x}\rightarrow[\sigma(y),x]$. In this case we set
$\ell_{B}(\Gamma,x,y;t):=\ell_{[\sigma(y),x]}((\Gamma_{\geq y})_{x};t)=\sum_{\sigma(y)\leq x^{\prime}\leq x}h((\Gamma_{\geq y})_{x^{\prime}};t)(-1)^{\rho_{B}(x^{\prime},x)}g([x^{\prime},x]^{*};t).$
We set $\ell_{B}(\Gamma,x,y;t)=0$ if $\sigma(y)\nleq x$.
@@
[KS16, Corollary 4.5] Let $\sigma:\Gamma\to B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ of rank $r$ and an Eulerian poset $B$. Then the local $h$-polynomial $\ell_{B}(\Gamma;t)$ has symmetric coefficients centered at $r/2$.
Remark 2.17.
Let $\sigma:\Gamma\rightarrow B$ be a strong formal subdivision between a lower Eulerian poset $\Gamma$ and an Eulerian poset $B$. We recover the $h$-polynomial of $\Gamma$ by the formula:
@@
Theorem 2.18.
(Katz-Stapledon decomposition of local $h$-polynomial)[KS16, Corollary 4.7] Let $\tau:\Omega\rightarrow\Gamma$ and let $\sigma:\Gamma\rightarrow B$ be strong formal subdivisions lower Eulerian posets $\Omega,\Gamma$ and Eulerian poset $B$. Then for $z\in\Omega$, $x\in\Gamma$,
@@
The boolean algebra on $r$ elements consists of all subsets of a set of cardinality $r$ and forms an Eulerian poset under inclusion of rank $r$. It is the face poset of an (r1)(r-1)-dimensional simplex. A poset $B$ with minimal element $\hat{0}$ is simplicial if for every $x$ in $B$, the interval $\hat{0},x]$ is a Boolean algebra.
@@
The local $h$-polynomial of strong formal subdivision is not necessarily non-negative (see [KS16, Example 5.6] and Corollary 3.12).
For a polyhedral subdivision $\mathcal{S}$ of a polytope $P$, we denote its local $h$-polynomial by $\ell_{P}(\mathcal{S};t)$, where the underlying strong formal subdivision is between the face poset of $\mathcal{S}$ and the face poset of $P$. Recall that $\sigma(F),F\in\mathcal{S}$ is defined as the minimal face of $P$ containing $F$. For any face $Q$ of $P$ containing $\sigma(F)$, we may consider the corresponding local $h$-polynomial:
$\ell_{P}(\mathcal{S},Q,F;t):=\ell_{[\emptyset,P]}(\mathcal{S},Q,F;t)=\ell_{[\sigma(F),Q]}(\operatorname{St}_{\mathcal{S}|_{Q}}(F);t).$
@@
[KS16, Theorem 6.1] Let $\mathcal{S}$ be a polyhedral subdivision of a polytope P. Then for every cell $F$ of $\mathcal{S}$ and face $Q$ of $P$ containing $F$, the local $h$-polynomial $\ell_{P}(\mathcal{S},Q,F;t)$ has non-negative symmetric coefficients. Moreover, if $\mathcal{S}$ is a regular subdivision, then the coefficients of $\ell_{P}(\mathcal{S},Q,F;t)$ are unimodal.
@@
In [KS16, Theorem 6.1] the condition ‘‘symmetric’’ is written only for regular subdivisions. But due to [KS16, Corollary 4.5] the local $h$-polynomial has symmetric coefficients for arbitrary strong formal subdivision, therefore we include the condition ‘‘symmetric’’ for arbitrary polyhedral subdivision. We also removed the word ‘‘rational’’ because the results from [Kar19] allows this (see e.g. [BKN24, Theorem 2.11]). The latter remark is non-trivial since there are non-rational convex polytopes which are not combinatorially equivalent to any rational polytopes; see e.g. [Zie08].
The local $h$-polynomials are non-negative in the polyhedral case because they appear when applying the Beilinson-Bernstein-Deligne decomposition theorem from [BBD81] to the toric morphisms (see [KS16, Lemma 6.4] and [CMM18]; see also [GKZ94, Notes and References, Part 2, Chapter 11, p. 514]).
2.3 $\mathbf{h^{*}}$- and local $\mathbf{h^{*}}$-polynomials of lattice polytopes
In this subsection we recall Ehrhart theory including (local) $h^{*}$-polynomials and Katz-Stapledon decomposition formulas for (local) $h^{*}$-polynomial from [KS16]. We recommend [BR07] as an introduction to Ehrhart theory including $h^{*}$-polynomials and [BKN24, §2] for a survey on (local) $h^{*}$-polynomials of polyhedral subdivisions and Katz-Stapledon decomposition formulas.
Let $P$ be a non-empty lattice polytope in a lattice $\mathbb{Z}^{d}$. Consider the function $f_{P}(m)=\#(mP\cap\mathbb{Z}^{d})$, for $m\in\mathbb{Z}_{>0}$. By Ehrhart’s theorem ([Ehr62] and e.g. [BR07, Section 3.4]), $f_{P}(m)$ is a polynomial of degree $\dim P$, called the Ehrhart polynomial of $P$, and $f_{P}(0)=1$. It follows that we can write